OpenStudy (austinl):

A point particle with charge q = 15 µC is placed at a corner of a cube of edge a = 5.6 cm. What is the flux through (a) each cube face forming that corner and (b) each of the other cube faces?

OpenStudy (austinl):

Here is my drawing.

OpenStudy (austinl):

(a) is zero, I am stuck on (b)

OpenStudy (austinl):

@UnkleRhaukus Are you familiar with this material. Electrostatics?

OpenStudy (austinl):

@skullpatrol Are you good with electrostatics?

OpenStudy (anonymous):

why is your cube having so many mini cubes in it? :D

OpenStudy (austinl):

Because, in the initial problem, the charge is on one corner of a cube, so you need to put more cubes around the original so that the charge is at the center of a larger cube.

OpenStudy (unklerhaukus):

this is just gauss's law,

OpenStudy (austinl):

I tried that, \(\Large{\Phi=\frac{q}{\epsilon_{0}}}\)

OpenStudy (unklerhaukus):

the electric flux goes out in straight lines from the point,

OpenStudy (anonymous):

ohhhww. that makes sense i guess wow.. yea.. so now you can use gauss's law and do it u are right.. the flux IS q/e_0 but that flux is passing through ALL those surface equally.. so its just a matter of dividing by the number of surfaces!

OpenStudy (unklerhaukus):

the flux through the adjoining surfaces will be zero , and the flux thought the other surface can be found from using symmetry,

OpenStudy (austinl):

Will it just be zero?

OpenStudy (anonymous):

no no.. its not zero.. think a little man.. u are on the right path!

OpenStudy (unklerhaukus):

the outer surfaces will get some flux , if you look at your diagram count all the outer surfaces , on the surrounding cubes , all the flux must go through these

OpenStudy (austinl):

4 cubes, with 6 sides each? And don't count those sides that are along the line of the flux.

OpenStudy (austinl):

So... flux through 20 different cube sides?

OpenStudy (austinl):

24 different sides*

OpenStudy (unklerhaukus):

there are more than 4 cubes, but yes 24 different sides,

OpenStudy (anonymous):

correct 24 sides!!

OpenStudy (anonymous):

i give u medal.. cause this was an awesome problem.. !!!!

OpenStudy (anonymous):

glad i saw this :D :D!!

OpenStudy (unklerhaukus):

so what is the total flux? and what portion of this flux goes thought each of those sides?

OpenStudy (austinl):

Okay, sorry about that. Having computer issues. \(\Large{\Phi_{total}=175,439}\) \(\Large{\Phi_{surfaces}=7,309.94}\) But that doesn't seem correct.

OpenStudy (anonymous):

are you sure u did that right? :P

OpenStudy (anonymous):

i get 70590 Nm^2/C as the answer!

OpenStudy (austinl):

For the first part I have, \(\Large{\Phi=\frac{1.5\times10^{-6}}{8.55\times10^{-12}}=175,439}\) \(\Large{\frac{\Phi}{24}=7309.94}\) Now, it says that it is wrong. What did I do wrong?

OpenStudy (anonymous):

epsilon zero = 8.854 * 10^-12 Farads/ meter!

OpenStudy (austinl):

*facepalm* Thank you.

OpenStudy (anonymous):

Also.. its 15 micro NOT 1.5 micro..

OpenStudy (anonymous):

don't *facepalm* u need ur glasses i guess :D :D jk!